# Generalization of Geometric Graph Neural Networks

### v1

@@ Abstract
In this paper, we study the generalization capabilities of geometric graph neural networks (GNNs). We consider GNNs over a geometric graph constructed from a finite set of randomly sampled points over an embedded manifold with topological information captured. We prove a generalization gap between the optimal empirical risk and the optimal statistical risk of this GNN, which decreases with the number of sampled points from the manifold and increases with the dimension of the underlying manifold. This generalization gap ensures that the GNN trained on a graph on a set of sampled points can be utilized to process other unseen graphs constructed from the same underlying manifold. The most important observation is that the generalization capability can be realized with one large graph instead of being limited to the size of the graph as in previous results. The generalization gap is derived based on the non-asymptotic convergence result of a GNN on the sampled graph to the underlying manifold neural networks (MNNs). We verify this theoretical result with experiments on both Arxiv dataset and Cora dataset.
@@ I Introduction
In [20], the authors prove generalization bound of GNNs with VC-dimension which is commonly used for convolutional neural networks. The authors in [21] analyze the generalization of a single-layer GNN based on stability analysis, which is further extended to a multi-layer GNN in [22]. In [23], the authors give a novel PAC-Bayesian analysis on the generalization bound of GNNs across arbitrary subgroups of training and testing datasets. The authors derive generalization bounds for GNNs via transductive uniform stability and transductive Rademacher complexity in [24, 25, 26]. The authors in [27] propose a size generalization analysis of GNNs correlated to the discrepancy between local distributions of graphs. The authors in [28] analyze the transductive generalization of a simple GNN with a contextual stochastic block model. We consider a continuous manifold model as a generative model for the graphs, which provides powerful theoretical support and is realistic to model real-world applications.
Generalization analysis has also been carried out on the graph-level when considering the graph classification problem. In [29], the authors provide a generalization bound based on PAC-Bayesian analysis with the bound depending on the maximum degree of the graph and the spectral norms. In [30], the authors provide a generalization bound on message passing networks comparable to Rademacher bounds in recurrent neural networks. The previous generalization gaps scale with the size of graph without capturing the underlying common structures of the graphs. A generalization analysis is carried out in [31, 32], where the graphs are on message-passing neural networks on graphs that are randomly sampled from a collection of template random graph models. We study the limit of graphs as a manifold, which is more realistic in practice and is more suitable to process high-dimensional data.
In [33], the authors also study the generalization analysis of GNNs from a manifold perspective both on node and graph levels. This work studies the generalization of GNNs to the MNNs directly, while our work focuses on the generalization capability when we only have access to sampled graphs from the manifold. Most importantly, only one graph is needed to train the GNN and this allows the generalization of all the other unseen sampled graphs from the manifold.
@@ II-A Graph Convolutions and Graph Neural Networks
for each layer $l=1,2\cdots,L$. The graph filter ${\mathbf{h}}_{\mathbf{G}}^{lpq}({\mathbf{S}})$ maps the $q$-th feature of layer $l-1$ to the $p$-th feature of layer $l$ as (5). We denote the GNN like (7) as a mapping $\bm{\Phi}({\mathbf{H}},{\mathbf{S}},{\mathbf{x}})$ for the ease of presentation, where ${\mathbf{H}}\in{\mathcal{H}}\subset{\mathbb{R}}^{P}$ denotes a set of the graph filter coefficients at all layers and ${\mathcal{H}}$ denotes the set of all possible parameter sets.
@@ II-B Manifold Convolutions and Manifold Neural Networks
We consider a $d$-dimensional compact, smooth and differentiable submanifold ${\mathcal{M}}$ embedded in ${\mathbb{R}}^{\mathsf{N}}$. The embedding induces a probability measure $\mu$ over the manifold with density function $\rho:{\mathcal{M}}\rightarrow(0,\infty)$, which is assumed to be bounded as $0<\rho_{min}\leq\rho(x)\leq\rho_{max}<\infty$ for each $x\in{\mathcal{M}}$. Manifold signals [13] are likewise defined as scalar functions $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$. We use $L^{2}({\mathcal{M}})$ to denote $L^{2}$ functions over ${\mathcal{M}}$ with respect to measure $\mu$. The inner product of signals $f,g\in L^{2}({\mathcal{M}})$ is defined as
@@
Consider the Laplace operator is self-adjoint and positive-semidefinite and the manifold ${\mathcal{M}}$ is compact, ${\mathcal{L}}_{\rho}$ has real, positive and discrete eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$, written as ${\mathcal{L}}_{\rho}\bm{\phi}_{i}=\lambda_{i}\bm{\phi}_{i}$ where $\bm{\phi}_{i}$ is the eigenfunction associated with eigenvalue $\lambda_{i}$, which can be seen as the osillation mode. The eigenvalues are ordered in increasing order as $0<\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\ldots$, and the eigenfunctions are orthonormal and form an eigenbasis of $L^{2}({\mathcal{M}})$. When mapping a manifold signal onto one of the eigenbasis $\bm{\phi}_{i}$, we have
@@ III Convergence and Generalization Analysis of GNNs via MNNs
$[{\mathbf{W}}_{N}]_{ij}=K_{\epsilon}(x_{N}^{i},x_{N}^{j})=\frac{\alpha_{d}}{(d+2)N\epsilon^{d+2}}\mathbbm{1}\left(\frac{\|x_{N}^{i}-x_{N}^{j}\|}{\epsilon}\right),$ (15)
where $\alpha_{d}$ is the volume of the $d$-dimensional Euclidean unit ball and $\mathbbm{1}$ stands for a characteristic function. Graph input and target output signals ${\mathbf{x}}_{N},{\mathbf{y}}_{N}\in{\mathbb{R}}^{N}$ are supported on sampled points $X_{N}$ belonging to $L^{2}(X_{N})$, whose values are sampled from manifold input signal $f$ and target signal $g$ respectively, written explicitly as
@@ III-A Convergence of GNNs to MNNs
The difference of the GNN $\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N},\cdot)$ on a graph ${\mathbf{G}}_{N}$ sampled from the manifold ${\mathcal{M}}$ and MNN $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},\cdot)$ converges to zero as $N$ goes to infinity.
@@ III-B Generalization of GNNs
Suppose a GNN is trained over the given graph ${\mathbf{G}}_{N}$ and graph signal ${\mathbf{x}}_{N}$. A positive loss function is denoted as $\ell(\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N})$ to measure the estimation performance. The training intends to minimize the empirical risk defined as
$P_{E}^{*}=\min\limits_{{\mathbf{H}}\in{\mathcal{H}}}R_{E}({\mathbf{H}}):=\ell(\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}).$ (20)
Given the fact that we only have access to a training set with limited training samples and not the underlying probability distribution $\mu$, nor the manifold ${\mathcal{M}}$, we can only aim to minimize the empirical risk in practice.
Substantially, the goal of the GNN is to minimize the statistical risk, i.e.
$P_{S}^{*}=\min\limits_{{\mathbf{H}}\in{\mathcal{H}}}R_{S}({\mathbf{H}}):=\mathbb{E}_{X_{N}\sim\mu^{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right].$ (21)
@@
which measures the difference between the optimal empirical risk and the optimal statistical risk of the GNN over the sampled graph from the manifold. Based on the convergence results that we have derived, we can bound the generalization gap in the following theorem.
@@
$GA={\mathcal{O}}\left(\left(\frac{\log\frac{N}{\delta}}{N}\right)^{\frac{1}{d+4}}+\left(\frac{\log N}{N}\right)^{\frac{1}{d+4}}\right).$ (23)
We observe that the generalization gap $GA$ defined as (22) decreases with the number of sampled points over the manifold $N$. Unlike previous results, we do not place any restrictions on the VC dimension or Rademacher complexities of the functions. We require that the graphs be sampled from the same underlying manifold. Furthermore, the generalization gap increases with the manifold dimension $d$. That is to say, for a higher manifold complexity, we need to sample more points to achieve a satisfying generalization gap.
@@
1Remark. This conclusion can also be extended to multi-layer and multi-feature neural network architectures, as the neural network is cascaded by layers of filters and nonlinearities. The generalization gap may propagate across layers which indicates the increasing of the generalization gap of multi-layer and multi-feature GNNs with the size of the architecture, which we will further verify in Section IV.
@@ IV Simulations
Figure 1: In the top row, we plot the accuracy in the training and testing sets as a function of the number of nodes in the training set for the Arxiv dataset. In the bottom row, we plot the accuracy difference and the best linear fit. For the linear approximation, we consider the points whose training accuracy is below $98\%$.
@@
Figure 2: In the top row, we plot the loss in the training and testing sets as a function of the number of nodes in the training set for the Arxiv dataset. In the bottom row, we plot the accuracy difference and the best linear fit. For the linear approximation, we consider the points whose training accuracy is below $98\%$.
@@
Figure 4: In the top row, we plot the accuracy/loss in the training and testing sets as a function of the number of nodes in the training set for the Cora dataset. In the bottom row, we plot the accuracy difference and the best linear fit.
@@
All experiments used a NVIDIA GeForce RTX 3090 GPU. We used SGD with a learning rate of $0.005$, no weight decay, and trained for $1000$ epochs. The reported value corresponds to the one obtained at the last iterate. In all cases we verified that the GNN converges to a stable value.
@@ IV-A Arxiv Dataset
Layers Hidden Units
Pearson Correlation
Accuracy / Loss
Arxiv Dataset
$1$ $64$ 0.991/0.980-0.991/-0.980- 0.991 / - 0.980
$1$ $128$ 0.992/0.980-0.992/-0.980- 0.992 / - 0.980
$1$ $256$ 0.994/0.985-0.994/-0.985- 0.994 / - 0.985
$1$ $512$ 0.993/0.981-0.993/-0.981- 0.993 / - 0.981
$2$ $64$ $/-0.997$ 0.997
$2$ $128$ 0.997/0.998-0.997/-0.998- 0.997 / - 0.998
$2$ $256$ 0.991/0.993-0.991/-0.993- 0.991 / - 0.993
$2$ $512$ 0.991/0.994-0.991/-0.994- 0.991 / - 0.994
$3$ $64$ 0.997/0.998-0.997/-0.998- 0.997 / - 0.998
$3$ $128$ 0.994/0.996-0.994/-0.996- 0.994 / - 0.996
$3$ $256$ 0.982/0.988-0.982/-0.988- 0.982 / - 0.988
$3$ $512$ 0.982/0.987-0.982/-0.987- 0.982 / - 0.987
$4$ $64$ 0.997/0.998-0.997/-0.998- 0.997 / - 0.998
$4$ $128$ 0.993/0.994-0.993/-0.994- 0.993 / - 0.994
$4$ $256$ 0.983/0.989-0.983/-0.989- 0.983 / - 0.989
$4$ $512$ 0.984/0.988-0.984/-0.988- 0.984 / - 0.988
Cora Dataset
$2$ $16$ 0.980/0.925-0.980/-0.925- 0.980 / - 0.925
$2$ $32$ 0.969/0.966-0.969/-0.966- 0.969 / - 0.966
$2$ $64$ 0.969/0.978-0.969/-0.978- 0.969 / - 0.978
$3$ $16$ 0.989/0.975-0.989/-0.975- 0.989 / - 0.975
$3$ $32$ 0.968/0.973-0.968/-0.973- 0.968 / - 0.973
$3$ $64$ 0.987/0.976-0.987/-0.976- 0.987 / - 0.976
TABLE I: Pearson Correlation Coefficients for the linear fit computed on Figures 1, 2, and 4 for the Arxiv and Cora datasets respectively. For the Arxiv dataset the coefficient was computed using the points corresponding to training accuracy below $98\%$. For the Cora dataset, all points were used.
Our main goal in this experiment is to validate the theoretical bound derived in Theorem 1. We evaluated this bound both for the accuracy (Figure 1) as well as the loss (Figure 2). In each plot’s upper row, we show the training and testing accuracy/loss achieved by the trained GNN. Intuitively, when the number of nodes is small, the GNN can overfit the training set,i.e. perfectly classify all the nodes in the training set. As the number of nodes increases, the GNN fails to overfit, and the training accuracy falls below $100\%$. In the lower, row of Figure 1 and Figure 2, we plot both the generalization gap (the difference between training and testing accuracy/loss) and the best linear fit for the points whose training accuracy is below $98\%$. Visually, we can see that the generalization gap follows a linear pattern in log scale. This is consistent with the theoretical result of Theorem 1. Formally, we computed the Pearson correlation coefficient which measures the linear relationships between the data. In Table I we show that the coefficient is always above $0.98$ both for accuracy as well as loss. This translates into a strong correlation, empirically validating the claim that the generalization gap follows a linear relationship in a logarithmic scale with respect to the number of nodes present in the empirical risk.

### v2

@@ Abstract
In this paper, we study the generalization capabilities of geometric graph neural networks (GNNs). We consider GNNs over a geometric graph constructed from a finite set of randomly sampled points over an embedded manifold with topological information captured. We prove a generalization gap between the optimal empirical risk and the optimal statistical risk of this GNN, which decreases with the number of sampled points from the manifold and increases with the dimension of the underlying manifold. This generalization gap ensures that the GNN trained on a graph on a set of sampled points can be utilized to process other unseen graphs constructed from the same underlying manifold. The most important observation is that the generalization capability can be realized with one large graph instead of being limited to the size of the graph as in previous results. The generalization gap is derived based on the non-asymptotic convergence result of a GNN on the sampled graph to the underlying manifold neural networks (MNNs). We verify this theoretical result with experiments on multiple real-world datasets.
@@ I Introduction
In [20], the authors prove the generalization bound of GNNs with VC-dimension which is commonly used for convolutional neural networks. The authors in [21] analyze the generalization of a single-layer GNN based on stability analysis, which is further extended to a multi-layer GNN in [22]. The authors in [23] extend this analysis to deeper GNNs and analyze the approximation theory of GNNs in [24]. In [25], the authors give a novel PAC-Bayesian analysis on the generalization bound of GNNs across arbitrary subgroups of training and testing datasets. The authors derive generalization bounds for GNNs via transductive uniform stability and transductive Rademacher complexity in [26, 27, 28]. The authors in [29] propose a size generalization analysis of GNNs correlated to the discrepancy between local distributions of graphs. The authors in [30] analyze the transductive generalization of a simple GNN with a contextual stochastic block model. The results of these works ignore the underlying graph structures, which leads to generalization bounds increasing with the number of nodes – a behavior contrary to the one experienced in practice. We consider a continuous manifold model as a generative model for the graphs, which provides powerful theoretical support and is realistic to model real-world applications.
Generalization analysis has also been carried out on the graph-level when considering the graph classification problem. In [31], the authors provide a generalization bound based on PAC-Bayesian analysis with the bound depending on the maximum degree of the graph and the spectral norms. In [32], the authors provide a generalization bound on message passing networks comparable to Rademacher bounds in recurrent neural networks. The previous generalization gaps again scale with the size of graphs without capturing the underlying common graph structures. A generalization analysis is carried out in [33, 34], where the graphs are on message-passing neural networks on graphs randomly sampled from a collection of template random graph models. We study the limit of graphs as a manifold, which is widely accepted in physical dynamics [35], and images (the so-called image manifold) [36, 37] to process high-dimensional data. In [38], the authors also study the generalization analysis of GNNs from a manifold perspective both on node and graph levels. This work studies the generalization of GNNs to the MNNs directly in a statistical way, while our work focuses on the generalization capability when we only have access to sampled graphs from the manifold. This provides a more direct and applicable framework for evaluating practical performances as we cannot access the continuous manifold model directly. Most importantly, only one graph is needed to train the GNN and this allows the generalization of all the other unseen sampled graphs from the manifold.
@@ II-A Graph Convolutions and Graph Neural Networks
for each layer $l=1,2\cdots,L$, with ${\mathbf{x}}_{l}^{p}\in{\mathbb{R}}^{N}$. The graph filter ${\mathbf{h}}_{\mathbf{G}}^{lpq}({\mathbf{S}})$ maps the $q$-th feature of layer $l-1$ to the $p$-th feature of layer $l$ as (5). We denote the GNN like (7) as a mapping $\bm{\Phi}({\mathbf{H}},{\mathbf{S}},{\mathbf{x}})$ for the ease of presentation, where ${\mathbf{H}}\in{\mathcal{H}}\subset{\mathbb{R}}^{P}$ denotes a set of the graph filter coefficients at all layers and ${\mathcal{H}}$ denotes the set of all possible parameter sets.
@@ II-B Manifold Convolutions and Manifold Neural Networks
We consider a $d$-dimensional compact, smooth and differentiable submanifold ${\mathcal{M}}$ embedded in ${\mathbb{R}}^{\mathsf{D}}$. The embedding induces a probability measure $\mu$ over the manifold with density function $\rho:{\mathcal{M}}\rightarrow(0,\infty)$, which is assumed to be bounded as $0<\rho_{min}\leq\rho(x)\leq\rho_{max}<\infty$ for each $x\in{\mathcal{M}}$. Manifold signals [13] are likewise defined as scalar functions $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$. We use $L^{2}({\mathcal{M}})$ to denote $L^{2}$ functions over ${\mathcal{M}}$ with respect to measure $\mu$. The inner product of signals $f,g\in L^{2}({\mathcal{M}})$ is defined as
@@
Consider the Laplace operator is self-adjoint and positive-semidefinite and the manifold ${\mathcal{M}}$ is compact, ${\mathcal{L}}_{\rho}$ has real, positive and discrete eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$, written as ${\mathcal{L}}_{\rho}\bm{\phi}_{i}=\lambda_{i}\bm{\phi}_{i}$ where $\bm{\phi}_{i}$ is the eigenfunction associated with eigenvalue $\lambda_{i}$, which can be seen as the osillation mode. The eigenvalues are ordered in increasing order as $0\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\ldots$, and the eigenfunctions are orthonormal and form an eigenbasis of $L^{2}({\mathcal{M}})$. When mapping a manifold signal onto one of the eigenbasis $\bm{\phi}_{i}$, we have
@@ III Convergence and Generalization Analysis of GNNs via MNNs
$[{\mathbf{W}}_{N}]_{ij}=K_{\epsilon}(x_{N}^{i},x_{N}^{j})=\frac{\alpha_{d}}{(d+2)N\epsilon^{d+2}}\mathbbm{1}_{[0,1]}\left(\frac{\|x_{N}^{i}-x_{N}^{j}\|}{\epsilon}\right),$ (15)
where $\alpha_{d}$ is the volume of the $d$-dimensional Euclidean unit ball and $\mathbbm{1}$ stands for a characteristic function, i.e. the function value is $1$ if and only if $\frac{\|x_{N}^{i}-x_{N}^{j}\|}{\epsilon}$ falls in $[0,1]$. We call graph ${\mathbf{G}}_{N}$ as a geometric graph as it contains the geometric information embedded in the underlying manifold model. Graph input and target output signals ${\mathbf{x}}_{N},{\mathbf{y}}_{N}\in{\mathbb{R}}^{N}$ are supported on sampled points $X_{N}$ belonging to $L^{2}(X_{N})$, whose values are sampled from manifold input signal $f$ and target signal $g$ respectively, written explicitly as
@@ III-A Convergence of GNNs to MNNs
The difference of the GNN $\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N},\cdot)$ on a graph ${\mathbf{G}}_{N}$ sampled from the manifold ${\mathcal{M}}$ and MNN $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},\cdot)$ converges to zero as $N$ goes to infinity when the input signal has large enough bandwidth.
@@ III-B Generalization of GNNs
Suppose a GNN is trained over the given graph ${\mathbf{G}}_{N}$ and graph signal ${\mathbf{x}}_{N}$. A positive loss function $\ell$ is used to measure the estimation performance. We assume that the loss function is normalized Lipschitz continuous defined as follows
Assumption 2.
(Normalized Lipschitz loss function) The loss function $\ell$ is normalized Lipschitz continuous, i.e., $|\ell({\mathbf{y}}_{i},{\mathbf{y}})-\ell({\mathbf{y}}_{j},{\mathbf{y}})|\leq\|{\mathbf{y}}_{i}-{\mathbf{y}}_{j}\|_{2}$, with $\ell({\mathbf{y}},{\mathbf{y}})=0$.
The training of GNNs intends to minimize the empirical risk $P_{E}^{*}=\min\limits_{{\mathbf{H}}\in{\mathcal{H}}}R_{E}({\mathbf{H}})$ with $R_{E}({\mathbf{H}})$ defined as
$R_{E}({\mathbf{H}}):=\ell(\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}).$ (20)
Given the fact that we only have access to a training set with limited training samples and not the underlying probability distribution $\mu$, nor the manifold ${\mathcal{M}}$, we can only aim to minimize the empirical risk in practice. Substantially, the goal of the GNN is to minimize the statistical risk $P_{S}^{*}=\min\limits_{{\mathbf{H}}\in{\mathcal{H}}}R_{S}({\mathbf{H}})$, with $R_{S}({\mathbf{H}})$ defined as
$R_{S}({\mathbf{H}}):=\mathbb{E}_{X_{N}\sim\mu^{N}}\left[\ell\left(\bm{\Phi}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{N}),{\mathbf{y}}_{N}\right)\right],$ (21)
@@
which measures the difference between the optimal empirical risk and the optimal statistical risk of the GNN over the sampled graph from the manifold. We note that by defining it in this way, the actual training loss, reflected by $P_{E}^{*}$, can be used to upper bound $P_{S}^{*}$. This provides a performance guarantee of the GNN over the whole dataset. Based on the convergence results that we have derived, we can bound the generalization gap in the following theorem.
@@
${GA={\mathcal{O}}\left(\left(\frac{\log\frac{N}{\delta}}{N}\right)^{\frac{1}{d+4}}\right).}$ (23)
We observe that the generalization gap $GA$ defined as (22) decreases with the number of sampled points over the manifold $N$. Unlike previous results, we do not place any restrictions on the VC dimension or Rademacher complexities of the functions. We require that the graphs be sampled from the same underlying manifold, . Furthermore, the generalization gap increases with the manifold dimension $d$. That is to say, for a higher manifold complexity, we need to sample more points to achieve a satisfying generalization gap. While in most practical scenarios, manifolds are often low-dimensional, especially those associated with physical systems or image manifolds, as demonstrated in [35, 37, 43].
@@
1Remark 1. This conclusion can also be extended to multi-layer and multi-feature neural network architectures, as the neural network is cascaded by layers of filters and nonlinearities. The generalization gap may propagate across layers which indicates the increasing of the generalization gap of multi-layer and multi-feature GNNs with the size of the architecture, which we will further verify in Section IV.
[…]
1Remark 2. The analysis of the generalization of GNNs on node classification tasks can be extended to graph classification tasks if we consider a linear combination of the node-wise outputs of the GNNs. The empirical risk $R_{E}$ and the statistical risk $R_{S}$ are defined as the summation of loss functions over a set of graphs, with each graph sampled from a manifold within a manifold set.
@@ IV Simulations
Figure 1: GNN with $3$ layers and varying numbers of hidden units $\{64,128,256,512\}$ with a varying number of nodes in the training set trained in the Arxiv dataset. In the top row, we plot the accuracy in the training and testing sets as a function of the number of nodes in the training set for the Arxiv dataset. In the bottom row, we plot the accuracy difference and the best linear fit. For the linear approximation, we consider the points whose training accuracy is below $98\%$.
@@
Figure 2: GNN with $3$ layers and varying numbers of hidden units $\{64,128,256,512\}$ with a varying number of nodes in the training set trained in the Arxiv dataset. In the top row, we plot the loss in the training and testing sets as a function of the number of nodes in the training set for the Arxiv dataset. In the bottom row, we plot the accuracy difference and the best linear fit. For the linear approximation, we consider the points whose training accuracy is below $98\%$.
@@
Figure 4: Cora dataset: GNN with $\{2,3\}$ layers and varying numbers of hidden units $\{16,32,64\}$ with a varying number of nodes in the training set trained in the Cora dataset. In the top row, we plot the accuracy/loss in the training and testing sets as a function of the number of nodes in the training set for the Cora dataset. In the bottom row, we plot the accuracy difference and the best linear fit.
(a) Two Layers
(b) Three Layers
(c) Two Layers
(d) Three Layers
Figure 5: CiteSeer dataset: GNN with $\{2,3\}$ layers and varying numbers of hidden units $\{16,32,64\}$ with a varying number of nodes in the training set trained in the CiteSeer dataset. In the top row, we plot the accuracy/loss in the training and testing sets as a function of the number of nodes in the training set for the Cora dataset. In the bottom row, we plot the accuracy difference and the best linear fit.
(a) Two Layers
(b) Three Layers
(c) Two Layers
(d) Three Layers
Figure 6: PubMed dataset: GNN with $\{2,3\}$ layers and varying numbers of hidden units $\{16,32,64\}$ with a varying number of nodes in the training set trained in the PubMed dataset. In the top row, we plot the accuracy/loss in the training and testing sets as a function of the number of nodes in the training set for the Cora dataset. In the bottom row, we plot the accuracy difference and the best linear fit.
@@
All experiments used a NVIDIA GeForce RTX 3090 GPU. We used SGD with a learning rate of $0.005$, no weight decay, and trained for $1000$ epochs. The reported value corresponds to the one obtained at the last iteration. In all cases we verified that the GNN converges to a stable value.
@@ IV-A Arxiv Dataset
Our main goal in this experiment is to validate the theoretical bound derived in Theorem 1. We evaluated this bound both for the accuracy (Figure 1) as well as the loss (Figure 2). In each plot’s upper row, we show the training and testing accuracy/loss achieved by the trained GNN. Intuitively, when the number of nodes is small, the GNN can overfit the training set,i.e. perfectly classify all the nodes in the training set. As the number of nodes increases, the GNN fails to overfit, and the training accuracy falls below $100\%$. In the lower, row of Figure 1 and Figure 2, we plot both the generalization gap (the difference between training and testing accuracy/loss) and the best linear fit for the points whose training accuracy is below $98\%$. Visually, we can see that the generalization gap follows a linear pattern in log scale. This is consistent with the theoretical result of Theorem 1. Formally, we computed the Pearson correlation coefficient which measures the linear relationships between the data. In Table LABEL:table:pearson we show that the coefficient is always above $0.98$ both for accuracy as well as loss. This translates into a strong correlation, empirically validating the claim that the generalization gap follows a linear relationship in a logarithmic scale with respect to the number of nodes present in the empirical risk.